3. A piano string of length 76cm and mass per unit length of 6.91 x 10-3
kg/m is under a
tension of value 40N. What is the frequency of a note sounded by this string
1
Expert's answer
2020-07-02T17:16:22-0400

Length of the piano string =76cm = 0.76 m


"\mu" = 6.91 x 10-3 kg/m


Tension of the value(T)= 40N


"f=[\frac{T}{\mu}]^{\frac{\frac{1}{2}}{2L}}"


"f=[\frac{40}{6.91 * 10^{-3}} ]^{\frac{\frac{1}{2}}{2*0.76}}"


"f = 50.1 Hz"